OpenStudy (nightmarenight):

Evaluate the function f(x) = 3x – 2, when x = –1. –9 –5 1 5

OpenStudy (a1234):

f(-1) = 3(-1) - 2 f(-1) = -3 - 2 f(-1) = -5

OpenStudy (the_fizicx99):

Simply plug in -1 for x and keep simplifying F(-1) = 3(-1) - 2 f(-1) = -3 - 2 f(-1) = -5